OpenStudy (anonymous):

Write and solve an equation for each of the following problems, then answer it. Ian has 10 fewer dimes than quarters. He has 5.30. Find the number of quarters and dimes Ian has.

OpenStudy (anonymous):

Let number of dimes be X Number of quarters = X + 10 Value of dime = (x)0.1 Value of quarters = (x+10)(0.25) Total value = 5.30 = 0.1x + (0.25)(x+10) = 0.1x + 0.25x + 2.5 = 0.35x + 2.5 = 5.3 x =?

OpenStudy (anonymous):

x=8. But why isn't it d=z-10 and q=z since it says it's that Ian has 10 fewer dimes than quarters?